# ScalingNoise: Scaling Inference-Time Search for Generating Infinite Videos

### v2

@@ 3.2 Formulation of Video Generation Inference
3.3 Reward Design
One-Step Denoising. Throughout our evaluation process, the actions in the MDP sample initial noises, which is typically Gaussian noise. Therefore, it is difficult to assess. A key challenge arises from the difficulty of directly assessing this initial noises. To address this, we propose a one-step evaluation approach. Specifically, we utilize the predicted $\text{predicted }{\hat{v}_{\tau_{0}}}$ component from Eq. 4 above as the target for evaluation. The detailed formulation is presented as:
${\hat{v}}_{\tau_{0}}=\frac{v_{\tau_{t}}-\sigma_{\tau_{t}}\bm{\epsilon}_{\theta}(v_{\tau_{t}},\tau_{t})}{\alpha_{\tau_{t}}}.$ (6)
Our evaluation method leverages a single DDIM step to accurately and efficiently assess the initial noise. In contrast to the original brute-force approach, which fully denoises the initial noises into a clear video, our technique is less computationally intensive, though it may yield suboptimal results. The brute-force method, while thorough, requires significant computational resources, rendering it impractical for scaling to long video generation.
Consistency Reward. After obtaining an evaluable object, we need to design a long-term reward to evaluate the overall consistency of the video and prevent the accumulation of errors. The reward function requires careful design. Specifically, we take into account not only the video clip currently being generated by the action but also the states of previous nodes. To this end, we select fully denoised, clear video frames $v_{a}$ from several frames prior as anchor points and assess the consistency of the video within the current window after one-step denoising. This approach substantially reduces inconsistencies in video generation. In our experiments, we employ a DINO [6] model and calculate the reward using the following formula:
$\Phi:=\frac{1}{2(T-1)}\sum_{i=1}^{M}\left(\langle d_{\text{a}}\cdot d_{i}\rangle+\langle d_{i}\cdot d_{i-1}\rangle\right),$ (7)
[…]
3.4 Action Design
Search Framework. Once equipped with a rewards model (Section 3.3), VDM can leverage any planning algorithm for search, as demonstrated in  [21]. We employ beam search (Fig. 1 (b)), a robust planning technique that efficiently navigates the search tree space, effectively balancing exploration and exploitation to identify high-reward trajectories. Specifically, beam search constructs a search tree iteratively, with each node representing a state and each edge denoting an action and the resulting state transition. To steer VDM toward the most promising nodes, the algorithm relies on a state-action reward function, $\Phi$ in Eq. 7. To promote diversity among response candidates at each step, we maintain $K$ distinct trajectories. From a tilted distribution, we sample $N$ initial noise instances, generating $K\times N$ candidates for the current step. The reward model evaluates each candidate, and the top-$K$ candidates with the highest scores are selected as responses for that step. This sampling and selection process iterates until the full response sequence is generated. Further details and the pseudo-code for this planning algorithm are provided in Algorithm 1.
$v_{i}=\bm{F}_{\text{low}}^{r}(v_{i})+\bm{F}_{\text{high}}^{r}(\eta)~{},$ (8)
where $\bm{F}$ denotes the FFT function, and $\eta$ is the Gaussian noise.
$\diamond$ A3 DDIM Inversion: Apply DDIM Inversion to re-noise the previous frame. It can be formulated as:
4 Experiment
In this section, we conduct experiments to answer the following two questions: 1. Does the long-term reward guided search yield higher-quality and consistent video compared with other inference-time scaling methods? 2. Does the one-step evaluation provide an efficient and accurate assessment of initial noises?
4.1 Baseline and Implementation details
[…]
Streamingt2v [24] 1 84.03 91.01 96.58 95.47 61.64 85.74
OpenSora v1.1 [32] 1 86.92 93.18 97.50 98.72 53.07 85.87
FreeNoise [53] 1 92.30 95.16 96.32 94.94 67.14 89.17
[…]
Vbench Dataset. Our approach is systematically evaluated using VBench [28], a video generation benchmark featuring 16 metrics crafted to thoroughly evaluate motion quality and semantic consistency. We select 40 representative prompts spanning all categories, generate 100 video frames, and analyze the model’s performance accordingly. We choose five VBench evaluation metrics for performance comparison: Subject Consistency, Background Consistency, Motion Smoothness, Temporal Flickering, and Imaging Quality. We maintain k distinct candidates in the beam and sample n completions for each one. Specifically, we set k = 2, 3 for beam size, and n = 5 with FIFO-Diffusion based on VideoCraft2 [10], n = 10 with ConsistI2V [55] chunk by chunk to balance the quality and efficiency. For the I2V model, we use the first image of the video generated by FIFO-Diffusion to guide the video generation. We consider two types of baselines: (1) Base Model: This serves as our baseline approach, employing a naive method that deliberately avoids any form of inference-time scaling. (2) Best of N (BoN): A widely adopted technique to improve model response quality during inference. For each prompt, we generate multiple responses, specifically 3 and 5 distinct outputs. We also select three state-of-the-art methods as baselines, namely StreamingT2V [24], Openasora v1.1 [32], and FreeNoise [53].
  • Inception Score (IS) [57] for frame-level quality and diversity.
We measure the two metrics using Latte [46] as a base model, which is a DiT-based video model trained on UCF-101 [64], employing FIFO-Diffusion as the paradigm for long video generation, configured with k = 2 and m = 5. We generate 2,048 videos with 128 frames each to calculate $\text{FVD}_{128}$, a version of FVD which uses 128-frames-long videos to compute the statistics, and randomly sample a 16-frame clip from each video to measure the IS score, following evaluation guidelines in StyleGAN-V [61]. As the base model, we choose StyleGAN-V, PVDM [101], FIFO-Diffusion, as they are representative open-sourced models.
Figure 3: (a) The two figures are boxplots showing the tendency of scaling beam sizes for ScalingNoise based on two paradigms, in order of FIFO-Diffusion and Chunk by chunk. (b) From Left to Right: Correction of reward model DINO and CLIP feature similarity score and final subject consistency. All points are generated by VideoCraft2.
4.2 Quantitative results
Scale Search Improves Video Consistency. We compare ScalingNoise with the baselines trained for long video generation in terms of multiple benchmarks. Vbench: As shown in Table 2, we find that the videos generated by ScalingNoise are significantly more preferred compared with the baseline. Videos generated by the base model decoding are of the lowest quality. While increasing inference compute via BoN shows improvement, they still fall short compared with ScalingNoise search. The computational cost of BoN(3) is comparable to that of our method when configured with k = 3. In contrast, the computational cost of BoN(5) significantly exceeds that of our approach. Although its consistency has improved, our method outperforms BoN(5) across all evaluated aspects. The long videos obtained using ScalingNoise search not only significantly ehnance consistency, but also provide high quality video frames. FVD and IS: As illustrated in Fig. 2, our approach outperforms all the compared methods including PVDM-L (400-400s) [101], which employs a chunked autoregressive generation strategy. Note that PVDM-L (400-400s) iteratively generates 16 frames conditioned on the previous outputs over 400 diffusion steps.
[…]
BoN 97.87 92.06 477.75
10-Step 97.14 91.59 79.67
ScalingNoise (Ours) 97.71 91.83 12.34
Benefits from Further Scaling Up Inference Compute. We next explore the effect of increasing inference-time computation on the response quality of the VDM at each step by varying the beam sizes. To ensure experimental fairness, we set n = 5 for FIFO-Diffusion, while for chunk-by-chunk methods, we maintain n = 10. This distinction arises because FIFO-Diffusion requires initializing only one initial noise per step, whereas the chunk-wise search process initializes one clip per step. Consequently, after a single DDIM denoising step, the chunk-wise approach needs to select the combination of multiple frames, resulting in a search space complexity significantly larger than that of FIFO-Diffusion. Setting n = 5 would yield an overly constrained search space and fail to align with the computational load of FIFO-Diffusion. We adopt the subject consistency from V-Bench, as mentioned earlier, as our evaluation criteria. We report the scores for long video generation achieved through ScalingNoise search, based on both FIFO-Diffusion and chunk-by-chunk paradigms, with beam sizes set to 1, 2, 3, and 4. The experimental results are illustrated in Fig. 3 (a). Since some prompts are static while others correspond to video actions with very large movements, this results in a significant variance. Our observations reveal that the performance of ScalingNoise, for both FIFO-Diffusion and chunk-wise strategies, improves steadily as the search beam size increases. This trend suggests that scaling inference-time computation effectively enhances the visual consistency capabilities of VDM.
One-Step Evaluation’s Efficiency and Accuracy. To evaluate computational efficiency and accuracy of our evaluation method, we examine the VBench evaluation metrics and inference time across different assessment methods. Using the VideoCraft2 [10] model (configured with DDIM=64), we generate videos with a fixed length, adopting 16 frames. We sample 10 candidate initial noises and employ our one-step evaluation method to select one. To ensure a fair comparison of initial noise quality, we fully denoise the selected noise and assess the consistency and quality of the resulting video. In addition, we test two baseline approaches: (1) BoN: selection after complete denoising for clarity, and (2) m-Step Evaluation: the initial noises are denoised for m steps via DDIM, followed by selection using the same reward model. Here we set m to 10. As shown in Table 3, our one-step evaluation method generates videos in just 12.34 seconds, enabling the assessment and selection of initial noises without compromising baseline performance. In contrast, the other two baselines require 79.67 seconds and 477.75 seconds, respectively. This efficiency allows for scalable search within the recent long video generation paradigm.
[…]
Local 92.16 66.83 88.94
Anchor 92.67 66.39 89.28
ScalingNoise (Ours) 93.14 67.91 89.60
4.3 Qualitative results
4.4 Ablation Study
In this section, we conduct ablation studies to evaluate the impact of each design component in ScalingNoise for long video generation. Unless otherwise specified, all experiments follow previous settings for fair comparison.
Figure 5: The upper part of this figure represents the original method, which involves a greedy approach to generating long videos. In contrast, the tree-structured searching process of ScalingNoise is outlined below: We highlight the path from the initial state to the best-performing generation, reporting the score for each node. Our prompt is “Red wine is poured into a glass. highly detailed, cinematic, arc shot, high contrast, soft lighting”.
As shown in Table 4, ScalingNoise achieves the best performance. As illustrated in Fig. 6, we present videos guided by different reward models. Our reward function not only considers the long-term consistency between the initial noise and anchor frame, but also effectively accounts for the cross-temporal influence of initial noise propagation across video frames within the denoising window.
Figure 6: Example Videos. Illustrations of long videos guided by different reward function. Row 1: Local reward only consider the quality of current denoised clip. Row 2: Anchor reward calculate the similarity of the anchor frame and the initial noise. Row 3: Ours combines the best of both, achieving long-term reward.
[…]
Different Reward Model. Then we investigated the effects of using different reward models (i.e., DINO [6] and CLIP [54]) to guide the search process. For each reward model, we generated a 16-frame video and, after one denoising step, scored it using DINO and CLIP, respectively. As shown in Fig. 3 (b), the vertical axis still represents the subject consistency score, while the horizontal axis represents the reward model scores. It can be observed that DINO’s similarity scores demonstrate a stronger alignment with the final video’s subject consistency compared to CLIP. Specifically, there exists a robust correlation between the consistency of the main subject in the final video and the feature space extracted by DINO, as measured by the cosine similarity of the video frames. In contrast to DINO, which effectively captures the features of the primary subject in each frame, CLIP tends to focus on extracting the overall features of the background. During video generation, inconsistencies predominantly stem from variations in the subject, while changes in the background remain relatively minor. Consequently, DINO provides a more accurate and reliable evaluation of subject consistency, making it a superior choice over CLIP for this purpose.
Effectiveness of Tilted Distribution. We investigate how each component of the tilted distribution impacts the quality of video generation. Table 5 summarizes the performance results for long video generation. We tested the performance of these sampling distributions separately, including (1) Random Distribution, the standard normal Gaussian distribution. (2) 2D FFT, which leverages frequency domain analysis to improve generation quality. (3) DDIM Inversion, a method that significantly enhances subject consistency across frames. (4) Inversion Resampling, an approach introduced to address specific limitations observed in earlier methods. The 2D FFT is an effective method for improving video quality. However, as the generated length increases, it can lead to a degradation in video quality. Although the DDIM reverse markedly enhances the subject consistency, it results in a significant reduction in the range of motion in the generated video. Therefore, we introduce Inversion Resampling to maintain the diversity. Integrating all methods into the base model yields a performance boost of +0.53%.
5 Conclusion & Limitation
6 VideoCrafter2
@@
(a) ”A lone figure in a hooded cloak stands atop a skyscraper, city lights sprawling below like glowing circuit board.”
@@
(b) ”A mystical, low-poly enchanted garden with glowing plants and a small pond where fireflies dance under a starry sky.”
@@
(c) ”A pair of tango dancers performing in Buenos Aires, 4K, high resolution.”
@@
(d) ”A tiger walks in the forest, photorealistic, 4k, high definition, 4k resolution.”
@@
(e) ”A tranquil, low-poly mountain village at dawn, with smoke rising from chimneys and birds flying over snowy peaks.”
@@
(f) ”Cinematic closeup and detailed portrait of a reindeer in a snowy forest at sunset.”
@@
(a) ”A dynamic, low-poly cityscape at night, with neon lights reflecting off wet streets and a lone cyclist riding through the rain.”
@@
(b) ”A bustling cityscape at sunset with skyscrapers reflecting golden light, people walking, and traffic moving swiftly.”
@@
(c) ”A peaceful, low-poly countryside with rolling hills, a windmill, and a farmer tending to his crops under a golden sunset.”
@@
(d) ”A horse race in full gallop, capturing the speed and excitement, 2K, photorealistic.”
@@
(e) ”A cozy, low-poly cabin in the woods surrounded by tall pine trees, with a warm light glowing from the windows and smoke curling from the chimney, 4k resolution.”
@@
(f) ”Impressionist style, a yellow rubber duck floating on the wave on the sunset, 4k resolution.”

### v3

@@ Appendix A Algorithm of ScalingNoise
Appendix B Baseline
Our approach is benchmarked against several methods:
  • FreeNoise [57]: We chose FreeNoise as a baseline because it is also a training-free method that can base the VideoCrafter2 [10] model, which also serves as our base model, to generate long videos. It employs a rescheduling technique for the initialization noise and incorporates Window-based Attention Fusion to generate longer videos.
  • Streaming T2V [24]: To assess our method’s effectiveness in generating longer videos, StreamingT2V was chosen as our baseline. Streaming T2V involves training a new model that uses an auto-regressive approach to produce long-form videos. Our prompt is "Red wine is poured into a glass. highly detailed, cinematic, arc shot, high contrast, soft lighting, 4k resolution. A spectacular fireworks display over Sydney Harbour, 4K, high resolution.".
  • OpenSora V1.1 [28]: a video diffusion model based on DiT [54], supports up to 120 frames, can generate videos at various resolutions, and has been specifically trained on longer video sequences to enhance its extended video generation capabilities.
Appendix C Benchmark
Vbench. Following is the detail of the five evaluation metrics in our paper: Subject Consistency assesses the uniformity and coherence of the primary subject across frames using DINO [6] features. Background Consistency is measured by the CLIP [59] feature similarity. Temporal Flickering [72] evaluates the frame-wise consistency and Motion Smoothness [38] assesses the fluidity and jittering of motion. Finally, we use MUSIQ [32] to predict the image quality which mainly considers the low-level distortions presented in the generated video frames.
Appendix D VideoCrafter2
@@
(a) "A lone figure in a hooded cloak stands atop a skyscraper, city lights sprawling below like glowing circuit board."
@@
(b) "A mystical, low-poly enchanted garden with glowing plants and a small pond
where fireflies dance under a starry sky."
@@
(c) "A pair of tango dancers performing in Buenos Aires, 4K, high resolution."
@@
(d) "A tiger walks in the forest, photorealistic, 4k, high definition, 4k resolution."
@@
(e) "A tranquil, low-poly mountain village at dawn,
with smoke rising from chimneys and birds flying over snowy peaks."
@@
(f) "Cinematic closeup and detailed portrait of a reindeer in a snowy forest at sunset."
@@
(a) "A dynamic, low-poly cityscape at night,
with neon lights reflecting off wet streets and a lone cyclist riding through the rain."
@@
(b) "A bustling cityscape at sunset with skyscrapers reflecting golden light,
people walking, and traffic moving swiftly."
@@
(c) "A peaceful, low-poly countryside with rolling hills, a windmill,
and a farmer tending to his crops under a golden sunset."
@@
(d) "A horse race in full gallop, capturing the speed and excitement, 2K, photorealistic."
@@
(e) "A cozy, low-poly cabin in the woods surrounded by tall pine trees, with a warm light
glowing from the windows and smoke curling from the chimney, 4k resolution."
@@
(f) "Impressionist style, a yellow rubber duck floating on the wave on the sunset, 4k resolution."